OpenStudy (anonymous):

Use the Racetrack Principle and the fact that sin 0 = 0 to show that sin x(less than or equal to) x for all x (greater than or equal to) 0.

OpenStudy (anonymous):

what the monkey is the "racetrack principle"?

OpenStudy (anonymous):

oh maybe that since the derivative of sine is cosine and cosine is bounded above by 1, then \[\frac{\sin(x)-\sin(0)}{x-0}\leq 1\] so \[\sin(x)\leq x\]

OpenStudy (anonymous):

ok I get that... is that still correct after looking at this pic